# MoFO: Momentum-Filtered Optimizer for Mitigating Forgetting in LLM Fine-Tuning

### v1

@@ Abstract
Recently, large language models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks. Typically, an LLM is pre-trained on large corpora and subsequently fine-tuned on task-specific datasets. However, during finetuning, LLMs may forget the knowledge acquired in the pretraining stage, leading to a decline in general capabilities. To address this issue, we propose a new fine-tuning algorithm termed Momentum-Filtered Optimizer (MoFO). The key idea of MoFO is to iteratively select and update the model parameters with the largest momentum magnitudes. Compared to full-parameter training, MoFO achieves similar fine-tuning performance while keeping parameters closer to the pre-trained model, thereby mitigating knowledge forgetting. Unlike most existing methods for forgetting mitigation, MoFO combines the following two advantages. First, MoFO does not require access to pre-training data. This makes MoFO particularly suitable for fine-tuning scenarios where pre-training data is unavailable, such as fine-tuning checkpoint-only open-source LLMs. Second, MoFO does not alter the original loss function. This could avoid impairing the model performance on the fine-tuning tasks. We validate MoFO through rigorous convergence analysis and extensive experiments, demonstrating its superiority over existing methods in mitigating forgetting and enhancing fine-tuning performance.
@@ 1 Introduction
In the literature, two classes of methods are commonly adopted to mitigate the forgetting: reply-based methods, and regularization-based methods.
  • Reply-based methods leverage pre-training data during the fine-tuning process (Rolnick et al., 2019; Wang et al., 2020; Ouyang et al., 2022). However, most open-source LLMs, such as the Llama series (Touvron et al., 2023), have not fully disclosed their pre-training datasets. Moreover, even with access to pre-training data, incorporating it into the fine-tuning process may significantly increase computational and memory costs.
@@
In this paper, we design a reply-free and regularization-free method to mitigate forgetting during the fine-tuning process. We propose the Momentum-Filtered Optimizer (MoFO). At each iteration, MoFO selects and updates only the parameters with the top $\alpha\%$ largest momentum magnitudes within each block, where $\alpha$ is the filtering hyperparameter. Here, the blocks refer to the parameters of different parts of the network (e.g., weight matrices and bias terms).
Our method is motivated by the following observation. The fine-tuning loss landscape of LLMs has many minima [^1] In this paper, we use the term ”minimum” (or ”minima” in the plural) to refer to a parameter configuration whose fine-tuning loss is near its lowest value in a small neighborhood, while acknowledging that this terminology may not strictly represent a local minimum of the fine-tuning loss function.). By conventional regularization-based methods, we can obtain a fine-tuned model that is close to the pre-trained model, thereby mitigating the forgetting effect. However, adding regularization will change the original loss function and degrade the performance of fine-tuning tasks. In contrast, fine-tuning with MoFO converges to a closer minimum of the original loss function, reducing the risk of forgetting without sacrificing the fine-tuning performance.
@@ 2.1 Motivation
It is hypothesized that the severity of forgetting during fine-tuning correlates with the distance between the fine-tuned model and the pre-trained model: in general, models that remain closer to the pre-trained model are less likely to experience forgetting (Kirkpatrick et al., 2017). This hypothesis motivates a line of regularized-based methods to mitigate forgetting (Kirkpatrick et al., 2017; Li et al., 2018; Zenke et al., 2017). However, adding regularization changes the original loss function, which can impair the model’s performance on the fine-tuning task (see evidence in Section 3). It is intriguing to design a new method to mitigate forgetting without changing the loss function. In this work, we propose a new optimization algorithm that can converge to minima (of the original fine-tuning loss) close to the pre-trained model without sacrificing the fine-tuning loss.
We first discuss how to keep the model closer to the pre-trained model. To achieve this goal, we recall the classical block coordinate descent (BCD) method (Tseng, 2001). We believe that BCD algorithm may converge to a point that is closer to the pre-trained model than full parameter fine-tuning. This is because BCD only updates a subset of parameters at each iteration. Compared to full parameter updates, this method usually involves smaller adjustments to the parameters.
@@ 2.3 Initial Exploration
We further verify whether the reduced parameter movement (by MoFO) effectively mitigates the forgetting of general capabilities acquired by pre-training. We evaluate the degree of forgetting from two perspectives: pre-training loss and evaluation benchmarks. First, we find MoFO has a lower pre-training loss compared to the Adam optimizer, indicating that MoFO memorizes pre-training data better. Second, in Table 1, we evaluate the accuracies of Pythia-160m, fine-tuned using both the Adam optimizer and MoFO, on widely-used common sense tasks, which measures the commonsense reasoning capabilities of LLMs. These tasks include HellaSwag (Zellers et al., 2019), ARC-Challenge, and ARC-Easy (Clark et al., 2018), which we will formally introduce in Section 3.1. The results show that MoFO achieves less accuracy degradations on all three tasks than the Adam optimizer. This indicates that MoFO not only helps maintain a smaller pre-training loss but also more effectively mitigates the forgetting of general capabilities.
Table 1: Pythia-160m’s performance on common sense tasks, after being fine-tuned with the Adam optimizer and MoFO. The results indicate that MoFO significantly mitigates catastrophic forgetting.
@@ 3.3 Continual Fine-Tuning
Table 4: The OP and BWT scores of TinyLlama-1.1B after fine-tuning on TRACE benchmark. Bold values denote the best results among these methods in each group.
@@ 5 Related Works
Researchers have proposed numerous methods to alleviate forgetting in continual learning, which involves learning a sequence of tasks. These methods are not limited to learning in a sequential manner and can be applied to broader paradigms. Generally, three primary approaches are used: reply-based methods, regularization-based methods, and architecture-based methods.
  • Reply-based methods leverage past experiences to facilitate the learning of new tasks. The most straightforward implementation of this approach is experience replay, which involves maintaining old samples in a buffer and replaying them during incremental training (Rolnick et al., 2019) with some variants enhancing performance or optimizing memory usage (Aljundi et al., 2019a; Hayes et al., 2019; Cha et al., 2021; Chaudhry et al., 2019b; Riemer et al., 2019b). Several other variants utilize gradient information from past tasks (Lopez-Paz and Ranzato, 2017; Riemer et al., 2019a; Chaudhry et al., 2019a; Farajtabar et al., 2020; Aljundi et al., 2019b; Chaudhry et al., 2021; Tiwari et al., 2022). In LLMs, several works (Yin et al., 2023; Wang et al., 2024; Ouyang et al., 2022) propose reply-based methods to mitigate forgetting. We emphasize that our MoFO method is orthogonal to reply-based methods. MoFO can be integrated into these replay strategies to further enhance their effectiveness.
  • Regularization-based methods introduce constraints to preserve old knowledge. Several studies add regularization terms to the loss functions to penalize parameter changes and mitigate forgetting (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Li et al., 2018; Ritter et al., 2018). Some other works apply regularization to the embedding or output changes (Li and Hoiem, 2017; Rannen et al., 2017; Buzzega et al., 2020; Huang et al., 2021). Unlike these approaches, MoFO does not modify the loss function and as a result, MoFO reached better SFT performance (see Table 2 and 3).
@@
In the realm of machine learning, BCD has also found applications (Nutini et al., 2022). For example, Luo et al. (2024) leverages BCD to perform memory-efficient fine-tuning of LLM and Xu and Zhang (2024) use random masking to perform this. A recent work, half fine-tuning approach (Hui et al., 2024) addresses catastrophic forgetting during fine-tuning of LLMs by selectively freezing half of the parameters during training. Our approach is akin to a more efficient greedy BCD, achieving superior performance in fine-tuning tasks and alleviating forgetting better.

### v2

@@ Abstract
Recently, large language models (LLMs) have demonstrated remarkable capabilities in a wide range of tasks. Typically, an LLM is pre-trained on large corpora and subsequently fine-tuned on task-specific datasets. However, during fine-tuning, LLMs may forget the knowledge acquired in the pre-training stage, leading to a decline in general capabilities. To address this issue, we propose a new fine-tuning algorithm termed Momentum-Filtered Optimizer (MoFO). The key idea of MoFO is to iteratively select and update the model parameters with the largest momentum magnitudes. Compared to full-parameter training, MoFO achieves similar fine-tuning performance while keeping parameters closer to the pre-trained model, thereby mitigating knowledge forgetting. Unlike most existing methods for forgetting mitigation, MoFO combines the following two advantages. First, MoFO does not require access to pre-training data. This makes MoFO particularly suitable for fine-tuning scenarios where pre-training data is unavailable, such as fine-tuning checkpoint-only open-source LLMs. Second, MoFO does not alter the original loss function. This could avoid impairing the model performance on the fine-tuning tasks. We validate MoFO through rigorous convergence analysis and extensive experiments, demonstrating its superiority over existing methods in mitigating forgetting and enhancing fine-tuning performance.
@@ 1 Introduction
In the literature, two classes of methods are commonly adopted to mitigate the forgetting: replay-based methods, and regularization-based methods.
  • Replay-based methods leverage pre-training data during the fine-tuning process (Rolnick et al., 2019; Wang et al., 2020; Ouyang et al., 2022). However, most open-source LLMs, such as the Llama series (Touvron et al., 2023), have not fully disclosed their pre-training datasets. Moreover, even with access to pre-training data, incorporating it into the fine-tuning process may significantly increase computational and memory costs.
@@
In this paper, we design a replay-free and regularization-free method to mitigate forgetting during the fine-tuning process. We propose the Momentum-Filtered Optimizer (MoFO). At each iteration, MoFO selects and updates only the parameters with the top $\alpha\%$ largest momentum magnitudes within each block, where $\alpha$ is the filtering hyperparameter. Here, the blocks refer to the parameters of different parts of the network (e.g., weight matrices and bias terms).
Our method is motivated by the following observation. The fine-tuning loss of LLMs has many minima [^1] In this paper, we use the term ”minimum” (or ”minima” in the plural) to refer to a parameter configuration whose fine-tuning loss is near its lowest value in a small neighborhood, while acknowledging that this terminology may not strictly represent a local minimum of the fine-tuning loss function. and these minima can vary significantly in their distances to the pretrained-model. We notice that the minima closer to the pre-trained model are less likely to experience forgetting. Inspired by this observation, MoFO manages to approach closer minima by dynamically updating a portion of parameters that are most effective for reducing fine-tuning loss at each iteration. Therefore, MoFO can reduce the risk of forgetting without sacrificing the fine-tuning performance.
@@ 2.1 Motivation
During fine-tuning, different training methods usually converge to different minima. We observe that these minima share similar fine-tuning loss but can vary significantly in their distances to the pre-trained model. Furthermore, minima closer are less likely to forget pre-training knowledge. Below, we provide an illustrative example of this phenomenon.
We conduct an experiment using the Pythia-160m model to illustrate this observation. We fine-tune this model on a subset of the FLAN dataset [^2] The subset we use is ‘definite_pronoun_resolution_10templates,’ available at https://huggingface.co/datasets/Muennighoff/flan. This is a preprocessed version of the FLAN dataset that incorporates updates made to the FLAN datasets since the release of the original FLAN. using two different optimizers: the Adam optimizer (Kingma and Ba, 2014) and the Lion optimizer (Chen et al., 2024). Figure LABEL:pythia_landscape_lion demonstrates that Adam and Lion converges to two different minima of the fine-tuning loss. Figure LABEL:pythia_landscape_lion(a) reveals that the two minima share similar fine-tuning loss. The minimum reached by Adam is significantly closer to the pre-trained model. Figure LABEL:pythia_landscape_lion(b) indicates that the minimum reached by Adam (the closer minimum) has a lower pre-training loss than that reached by Lion (the farther minimum). Furthermore, we evaluate the forgetting of pre-training knowledge in some common sense reasoning tasks, which include HellaSwag (Zellers et al., 2019), ARC-Challenge, and ARC-Easy (Clark et al., 2018). Table 1 shows that Adam suffers from less accuracy degradation on average, indicating better preservation of the pre-training knowledge.
Therefore, we conjecture that the severity of forgetting during fine-tuning correlates with the distance between the fine-tuned model and the pre-trained model: in general, models that remain closer to the pre-trained model are less likely to experience forgetting. Consequently, we consider whether it is possible to find a method superior to the Adam optimizer that not only reaches the minimum during fine-tuning but also converges closer to the pre-trained model, thereby mitigating forgetting more effectively.
Table 1: Pythia-160m’s accuracies on common sense tasks, after being fine-tuned with the Adam optimizer and Lion optimizer. Adam achieves less forgetting than Lion on average.
HellaSwag ARC-easy ARC-challenge Average
Pythia-160m 30.1 39.6 23.8 31.2
Adam 28.3 37.4 22.1 29.3
Lion 26.5 29.0 24.1 26.5
We first discuss how to keep the model closer to the pre-trained model. To achieve this goal, we recall the classical block coordinate descent (BCD) method (Tseng, 2001). We believe that the BCD algorithm may converge to a minimum that is closer to the pre-trained model than the default full parameter fine-tuning with Adam. This is because BCD only updates a subset of parameters at each iteration. Compared to full parameter updates, this method usually involves smaller adjustments to the parameters.
@@ 2.3 Initial Exploration
Table 2: Pythia-160m’s performance on common sense tasks, after being fine-tuned with the Adam optimizer and MoFO. The results indicate that MoFO significantly mitigates catastrophic forgetting. Bold values denote the best results among these optimizers.
@@
We further verify whether the reduced parameter movement (by MoFO) effectively mitigates the forgetting of general capabilities acquired by pre-training. We evaluate the degree of forgetting from two perspectives: pre-training loss and evaluation benchmarks. First, in Figure LABEL:pythia_landscape(b), we find MoFO has a lower pre-training loss compared to the Adam optimizer, indicating that MoFO memorizes pre-training data better. Second, in Table 2, we evaluate the accuracies of Pythia-160m, fine-tuned using both the Adam optimizer and MoFO, on some widely-used common sense tasks, which measures the commonsense reasoning capabilities of LLMs. The results show that MoFO achieves less accuracy degradations on all three tasks than the Adam optimizer. This indicates that MoFO not only helps maintain a smaller pre-training loss but also more effectively mitigates the forgetting of general capabilities.
@@ 3.3 Continual Fine-Tuning
Table 5: The OP and BWT scores of TinyLlama-1.1B after fine-tuning on TRACE benchmark. The results show that MoFO outperforms Full FT and HFT in continual learning and can combine well with traditional continual learning methods. Bold values denote the best results among these methods in each group.
@@ 5 Related Works
Researchers have proposed numerous methods to alleviate forgetting in continual learning, which involves learning a sequence of tasks. These methods are not limited to learning in a sequential manner and can be applied to broader paradigms. Generally, three primary approaches are used: replay-based methods, regularization-based methods, and architecture-based methods.
  • Replay-based methods leverage past experiences to facilitate the learning of new tasks. The most straightforward implementation of this approach is experience replay, which involves maintaining old samples in a buffer and replaying them during incremental training (Rolnick et al., 2019) with some variants enhancing performance or optimizing memory usage (Aljundi et al., 2019a; Hayes et al., 2019; Cha et al., 2021; Chaudhry et al., 2019b; Riemer et al., 2019b). Several other variants utilize gradient information from past tasks (Lopez-Paz and Ranzato, 2017; Riemer et al., 2019a; Chaudhry et al., 2019a; Farajtabar et al., 2020; Aljundi et al., 2019b; Chaudhry et al., 2021; Tiwari et al., 2022). In LLMs, several works (Yin et al., 2023; Wang et al., 2024; Ouyang et al., 2022) propose replay-based methods to mitigate forgetting. We emphasize that our MoFO method is orthogonal to replay-based methods. MoFO can be integrated into these replay strategies to further enhance their effectiveness.
  • Regularization-based methods introduce constraints to preserve old knowledge. Several studies add regularization terms to the loss functions to penalize parameter changes and mitigate forgetting (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Li et al., 2018; Ritter et al., 2018; Kumar et al., 2023). Some other works apply regularization to the embedding or output changes (Li and Hoiem, 2017; Rannen et al., 2017; Buzzega et al., 2020; Huang et al., 2021). Unlike these approaches, MoFO does not modify the loss function and as a result, MoFO reached better fine-tuning performance (see Table 3 and 4).
@@
In the realm of machine learning, BCD has also found applications (Nutini et al., 2022). For example, Luo et al. (2024) leverages BCD to perform memory-efficient fine-tuning of LLM and Xu and Zhang (2024) use random masking to perform this. In federated learning, Rothchild et al. (2020) adopts top-$k$ momentum value unsketch rather than our top-$k$ momentum filtering to tackle communication bottleneck and convergence issues. A recent work, Hui et al. (2024) addresses catastrophic forgetting during fine-tuning of LLMs by selectively freezing half of the parameters during training. Our approach is akin to a more efficient greedy BCD, achieving superior performance in fine-tuning tasks and alleviating forgetting better.